# HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models

### 11 Dec 2023 v1

@@ 3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the first problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [xu2023interdiff] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Fig. 4, the input to this model includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
Affordance estimation in 3D point clouds itself is a notably challenging problem [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive], especially in the context of 3D HOI generation involving textual prompt. To address the limited availability of training data, we propose a simplified approach that diverges from predicting point-wise contact labels. Our method is grounded in two key hypotheses. Firstly, we posit that interactions with an object involve a maximum of two contact points. Secondly, we identify eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the primary interacting parts in most HOI scenarios. These hypotheses streamline the complexity of the problem, making it more tractable without significantly compromising accuracy. For instance, this approach effectively models common interactions such as grasping an object with both hands, sitting actions involving the pelvis, or lifting with a single hand. Our empirical observations suggest that these simplifications are valid in a wide range of HOI applications, providing a balance between model complexity and real-world applicability.
@@
Specifically, at each diffusion time step $n$ of APDMWe note that APDM and HOI-DM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy affordance data ${\bm{y}}_{n}$ includes human contact labels $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, two object contact positions $\bm{y}_{n}^{o}\in\mathbb{R}^{2\times 3}$. In addition, our APDM also generates a binary object state label ${\bm{y}}_{n}^{s}\in\{0,1\}$ indicating whether it remains static or moves during the interaction. The generated binary label can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. APDM works similar to the diffusion denoising process described in Eq.(1). Its output include clean affordance information ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
@@ 3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [dhariwal2021diffusion] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Algorithm 1 Affordance-guided Interaction Correction
1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$
3:$K$ = 1
4:for all $t$ from $T$ to $1$ do
5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$
6:     Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$
7:     if $t=1$ then
8:         $K=100$
9:     end if
10:     for all $k$ from $K$ to $1$ do
11:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$
12:         ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$
13:     end for
14:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$
15:     $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$
16:end for
17:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can purturb the generated human and object motion at each diffusion step $t$ as in [xie2023omnicontrol, karunratanakul2023guided, kulkarni2023nifty],
@@
$G_{con}=\sum_{i\in\{1,2\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively. If there is just a single predicted contacting joint, $i=1$. $R(\cdot)$ converts the human joint’s local positions to global absolute locations, and $V(\cdot)$ obtains the object’s contact point sequence from the predicted mean of object pose ${\bm{\mu}}^{o}_{t}$.
Second, the generated motion of dynamic objects typically follows human movement. However, we observe that when the human interacts with a static object, such as sitting on a chair, the object appears slightly moved. To address this, we utilize the predicted object states ${\bm{y}}_{0}^{s}$ to immobilize the object’s movement in the generated samples if the state is static, ensuring that proper contact is established between the human and the static object. The objective is defined as
$\small{G_{sta}=\sum_{l=1}^{L}{\bm{y}}_{0}^{s}(l)\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
@@
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to space limit, we explain it in the supplementary material.
@@ 4 Experiments
Table 1: Quantitative results on the BEHAVE [bhatnagar22behave] dataset. We compare our method with three baselines adapted from existing models and the real data. The right arrow normal-$\rightarrow$ means closer to real data is better.
@@
MDM* [tevet2023human] 3.169 0.206 7.606 0.448 0.184 1.20%
PriorMDM* [shafir2023human] 4.352 0.215 8.186 0.416 0.270 21.42%
Ours (Full) 1.632 0.222 7.596 0.363 0.190 77.38%
@@
Supplementary Material

### 15 Mar 2024 v2

@@ 3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [54] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Fig. 4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
Affordance estimation in 3D point clouds itself is a notably challenging problem [30, 7, 22, 16, 29, 20, 21], especially in the context of 3D HOI generation involving textual prompt. In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios. It can effectively model common interactions such as grasping an object with both hands, sitting actions involving the pelvis and back, or lifting with a single hand. We use binary contact labels to determine which joints are in contact with the object. Subsequently, we predict eight corresponding contact points on the object surface, identified as the points closest to the selected body joints. Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
@@
Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and HOI-DM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$. The model is designed to predict both contact probabilities and contact positions. Subsequently, dynamic selection of contacting body joints is performed by considering predicted probabilities over a specific threshold $\tau$ (set to be 0.6). The corresponding contact points on the object are then determined based on the selected joints. APDM works similar to the diffusion denoising process described in Eq.(1). Besides, we utilize a large language model (ChatGPT) to determine whether the object state ${\bm{y}}_{0}^{s}\in\{0,1\}$ should be set to static (${\bm{y}}_{0}^{s}=1$) based on the textual description, which can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
@@ 3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [8] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Algorithm 1 Affordance-guided Interaction Correction 1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$. 2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$ 3:$K$ = 1 4:for all $t$ from $T$ to $1$ do 5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$ 6:     if $t=1$ then 7:         $K=100$ 8:     end if 9:     for all $k$ from $K$ to $1$ do # Separately perturb 10:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$ 11:     end for 12:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$ 13:end for 14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [53, 19, 23],
@@
$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th available contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively. $R(\cdot)$ converts the human joint’s local positions to global absolute locations, and $V(\cdot)$ obtains the object’s contact point sequence from the predicted mean of object pose ${\bm{\mu}}^{o}_{t}$.
Second, the generated motion of dynamic objects typically follows human movement. However, we observe that when the human interacts with a static object, such as sitting on a chair, the object appears slightly moved. To address this, we immobilize the object’s movement in the generated samples if the state is static (${\bm{y}}_{0}^{s}=1$), ensuring that proper contact is established between the human and the static object. The objective is defined as
$\small{G_{sta}={\bm{y}}_{0}^{s}\cdot\sum_{l=1}^{L}\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
@@
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in the supplementary material.
@@ 4 Experiments
4.1 Setup
Dataset. Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [3]. We then use ChatGPT to assist in generating three text descriptions for each HOI sequence, increasing the diversity of the data. Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects. It provides the human SMPL-H representation [27], the object mesh, as well as its 6DoF pose information in each HOI sequence. To ensure consistency in our approach, we follow the processing method used in HumanML3D [11] to extract representations for 22 body joints. All the models are trained to generate $L=196$ frames in our experiments. In the end, we have 1451 3D HOI sequences along with textual descriptions to train and evaluate our proposed approach. We follow the official train/test split on BEHAVE. We provide more details of the dataset, our annotation process, and annotated textual examples in the supplementary material.
In addition, we evaluate our approach on OMOMO dataset [25]. OMOMO focuses on full-body manipulation with hands. It consists of human-object interaction motion for 15 objects in daily life, with a total duration of approximately 10 hours. It provides text descriptions for each interaction motion. We utilize their object split strategy for both training and evaluation, ensuring the objects between the training and testing sets are different. Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset. More details of the OMOMO dataset are in the supplementary material.
Evaluation metrics. We first assess different models for human motion generation using standard metrics as introduced by [11], namely Fréchet Inception Distance (FID), R-Precision, and Diversity. FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder. R-Precision gauges the relevance between generated motions and their corresponding text prompts. Diversity evaluates the range of variation in the generated motions. Additionally, we compute the Foot Skating Ratio to measure the proportion of frames exhibiting foot skid over a threshold (2.5 cm) during ground contact (foot height $<$ 5 cm).
To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points. Ideally, we should develop similar metrics, e.g., FID, to evaluate the stochastic HOI generation. However, due to the limited data available in BEHAVE [3], training a motion encoder would produce biased evaluation results. To mitigate this issue, we resort to user studies to quantify the effectiveness of different models. Details will be introduced later.
4.2 Comparisons with Existing Methods
[…]
@@ 4.2 Comparisons with Existing Methods
MDM* [45] 3.169 0.206 7.606 0.448 0.184 0.87%
PriorMDM* [41] 4.352 0.215 8.186 0.416 0.270 16.53%
InterDiff [54] 7.192 0.170 7.686 0.506 0.218 0.00%
Ours (Full) 1.501 0.235 7.391 0.347 0.182 82.60%
Table 2: Quantitative results on OMOMO dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Real 1.539 0.193 4.833 - -
MDM* 9.155 0.137 5.353 0.768 0.191
PriorMDM 11.192 0.128 7.404 0.523 0.344
InterDiff 10.567 0.110 6.493 0.906 0.239
Ours 8.661 0.168 4.780 0.326 0.141
Figure 5: Qualitative comparisons of our approach and baseline methods on BEHAVE dataset. The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4. This contrasts with the baselines, which fail to achieve a similar level of realism and contact plausibility in the interactions. As an additional visual aid, the mesh color gradually darkens over time to represent progression. (Best viewed in color.)
Quantitative Results. Sec. 4.2 reports the quantitative results on BEHAVE dataset [3]. Compared with the baseline methods, our full method achieves the best performance. Specifically, it achieves state-of-the-art results in both FID, R-precision, and Diversity, underscoring its ability to generate high-quality human motions in the context of coherently interacting with objects. The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects.
We also show the results of our User Study in Tab. 4.2 to illustrate user preferences between our method and other baselines. In our user study, we sample 24 sequences from each method and 23 participants are asked to choose their most preferred generation results from these samples. This user study is designed to directly compare user preferences across different methods. The result indicates a strong preference for our method: it is favored over the baselines in a substantial majority of 82.60% of the cases.
Tab. 4.2 presents the quantitative results on the OMOMO dataset. Our method consistently outperforms other baselines by a considerable margin across all metrics. Notably, due to the distinctiveness of objects in the training and testing sets, the results in Tab. 4.2 indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
@@
FID $\downarrow$
R-precision $\uparrow$
User Study $\uparrow$
w/o Interaction Correction
Ours w/o CM 2.825 0.201 7.850 0.524 0.265 0.00%
Ours w/o pretrain 2.414 0.226 7.915 0.402 0.158 3.90%
[…]
Ours 1.808 0.217 7.718 0.416 0.205 11.69%
w/ Interaction Correction
Ours w/o ${M}^{o}$ CM 3.206 0.205 7.909 0.365 0.310 0.00%
Ours ${}^{Joint}$ 3.412 0.195 7.878 0.421 0.342 -
Ours w/o $G_{con}$ 1.983 0.205 7.826 0.417 0.196 -
Ours w/o $G_{sta}$ 1.636 0.213 7.785 0.367 0.181 -
Ours w/o $G_{smo}$ 1.665 0.205 7.840 0.370 0.182 -
Ours (Full) 1.501 0.235 7.391 0.347 0.182 84.41%
Figure 6: Visual results of different variants of our model in ablation studies on the BEHAVE dataset [3].
Qualitative Results. We showcase qualitative comparisons, rendered with SMPL [27] shapes, between our approach and the baseline methods in Fig. 5. It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g., the toolbox in Fig. 5 (b)). Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g., the chair in Fig. 5 (e)). In stark contrast, our approach excels in creating visually appealing and realistic HOIs. Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g., raising the trash bin with the right hand in Fig. 5 (a)).